# A Refreshed Similarity-based Upsampler for Direct High-Ratio Feature Upsampling

### v1

@@ 5.3.1 Model Visualization
Table V: Ablation study on the query-key feature alignment in Sec. 4.1 with the similarity measurement $\text{sim}(\cdot,\cdot)$ implemented by the PCDC block in Sec. 4.2 and the neighbor selected via the FNS strategy in Sec. 4.3.
@@ 5.3.2 Ablation Study
Ablation Study on Similarity Measurement. To comprehensively validate the effectiveness of our proposed PCDC-Block, we compare three different similarity measurements $\text{sim}(\cdot,\cdot)$, including (a) traditional inner product in Eq. (3); (b) replacing the PCDC layer in Fig. 5 with a vanilla convolution by inputting the concatenation of $\bar{\bm{q}}$ and $\bar{\bm{k}}$; (c) PCDC-Block in Fig. 5. From the results in Table VI, it is observed that the proposed PCDC-Block has the better capability to capture the relations between query-key pairs for precise similarity calculation and thus promotes the segmentation accuracy. Considering the evident gains over (a), the slight computational cost brought by our method is acceptable. Please refer to supplementary material for visual results.
@@
Ablation Study on Neighbor Selection. Table VII reports the effect of FNS on the segmentation performance achieved by our ReSFU. As seen, without adopting FNS, the model would suffer from a large performance drop, especially for SegFormer with a higher-ratio (×$\times 8$) upsampling process as presented in Fig. 12. For better clarity, we provide visual comparisons of upsampled features extracted by these two different network structures. From Fig. 17, we can clearly find that for different levels of features, i.e., $\bm{x}$ in Segmenter, $C2$, $C3$, and $C4$ in SegFormer, their upsampled feature maps without FNS always suffer from severe mosaic effects, especially for larger upsampling ratios. Fortunately, this unfavorable phenomenon can be effectively eliminated by adopting our proposed FNS strategy. It is worth mentioning that as explained in Sec. 4.3, FNS is directly implemented based on bilinearly-upsampled features, and it introduces no extra parameters and computational costs.
Table IX: Comparisons on FLOPS and the number of parameters, where FLOPs are tested with images of size 512$\times$512.
Method Segmenter-ViT-S Method Segmenter-ViT-S
FLOPs Params FLOPs Params
Bilinear 32.77G 22.04M FADE [12] 33.92G 22.14M
Deconv 35.16G 24.40M SAPA [2] 33.28G 22.09M
PixelShuffle [8] 35.16G 24.41M DySample [13] 32.90G 22.09M
Stack-JBU [22] 35.32G 22.20M ReSFU 33.88G 22.11M
CARAFE [7] 32.96G 22.21M / / /
@@
Under the default configurations with $K=3$, $D=32$, and $G=4$, the overhead of our ReSFU is comparable or even lower than other baselines (see Table IX). Taking into account performance, universality, and ease of deployment, our method has a better potential for real applications.
@@ 5.4.1 Medical Image Segmentation
Performance Comparison. The experiments are executed based on two widely-adopted datasets, i.e., Synapse for multi-organ segmentation, [^3] https://www.synapse.org/#!Synapse:syn3193805/wiki/217789 and the automated cardiac diagnosis challenge (ACDC). [^4] https://www.creatis.insa-lyon.fr/Challenge/acdc/ Specifically, Synapse is annotated with eight abdominal organs, i.e., aorta, gallbladder, spleen, left kidney, right kidney, liver, pancreas, spleen, and stomach. Following [48], we employ 18 computed tomography (CT) scans with 2,212 axial slices for training and 12 CT scans for validation. For ACDC, the involved magnetic resonance imaging (MRI) scans are annotated with left ventricle (LV), right ventricle (RV), and myocardium (MYO). Consistent to [48], we utilize 70 scans with 1,930 axial slices for training and 40 scans for validation. Quantitative evaluation is conducted based on the two typical metrics, including Dice (the higher the better) and the 95-th percentile of the Hausdorff Distance (HD95, the lower the better).
@@ 5.4.2 Instance Segmentation and Panoptic Segmentation
Specifically, for instance segmentation, we employ the classic Mask R-CNN [42] with ResNet50 as the backbone and replace the upsampling stages in FPN with ReSFU, as shown in Fig. 15 (b). All the results are quantitatively evaluated based on metrics of the average precision (AP) series, including mask AP, AP50, AP75, APS, APM, and APL. For panoptic segmentation, following [13], we choose Panoptic FPN [6] with ResNet50 as the backbone and adjust the three upsamplers in FPN, as presented in Fig. 15 (b). We report panoptic quality (PQ), PQ on things (PQth), PQ on stuff (PQst), segmentation quality (SQ), and recognition quality (RQ) [6]. The MS COCO dataset [67] is used for training and evaluating the two tasks.
@@
Figure 22: Visual comparison of upsampled features $\bm{\tilde{x}}$ obtained in an iterative upsampling (Fig. 9 (a)) and a direct ×$\times 4$ upsampling manners (Fig. 9 (b)) based on PSPNet-ResNet101.
@@ 5.5 More Discussions on Direct/Iterative Upsampling
Taking PSPNet as an example, for the comparing methods, e.g., CARAFE, FADE, SAPA, DySample, and ReSFU, we visually compare their upsampled features $\bm{\tilde{x}}$ obtained in the iterative ×$\times 2$ upsampling manner (Fig. 9 (a)) and the direct ×$\times 4$ upsampling manner (Fig. 9 (b)). As displayed in Fig. 22, with the high-ratio direct upsampling, CARAFE, FADE, SAPA, and DySample generate varying degrees of artifacts with more mosaics or blurrier boundaries compared with the iterative upsampling manner. Therefore, in all the experimental comparisons, we follow the default settings in the existing upsamplers and employ them in an iterative manner for better performance. Besides, we also observe that the proposed ReSFU achieves better visual effects with richer semantics and clearer boundaries under both upsampling manners. Therefore, for ease of deployment, we select the direct high-ratio upsampling manner as the default setting for ReSFU in experiments, which can eliminate the requirement of iterative guidance features.
@@ 6 Conclusion and Future Work
In this paper, for the fundamental upsampling design included in almost all current network architectures, we carefully reformulated the current similarity-based feature upsampling framework and thoroughly analyzed the limitations in methodological designs together with experimental visualizations. For each component involved in this pipeline, we meticulously proposed specific optimization designs and correspondingly constructed a refreshed similarity-based feature upsampling framework, called ReSFU. Through comprehensive model verification and ablation studies, we fully validated the role of each component and clearly shown the working mechanism underlying our proposed ReSFU. Based on different types of network architectures, extensive experiments substantiated the superiority of our proposed ReSFU beyond the existing baselines and demonstrated that our ReSFU can be adapted not only to different upsampling structures with flexible guidance manners but also to different applications, including medical image segmentation, instance segmentation, and panoptic segmentation.
Thanks to the favorable generality of our proposed ReSFU, apart from segmentation tasks, there is great potential to extend it to more applications, such as, the pansharpening task [68] or the multispectral and hyperspectral image fusion task [69] with the available HR input images as guidance. However, for the applications without HR reference images, e.g., image super-resoultion [70] and image generation [71], these guidance-based feature upsampling methods, e.g., FADE, SAPA, and our ReSFU, are difficult to apply. This is beyond our focus in this paper and deserves further investigation in our future work.

### v2

@@ 5.3.1 Model Visualization
Table V: Ablation study on the query-key feature alignment in Sec. 4.1 with the similarity measurement $\text{sim}(\cdot,\cdot)$ implemented by the PCDC-Block in Sec. 4.2 and the neighbor selected via the FNS strategy in Sec. 4.3.
@@ 5.3.2 Ablation Study
Ablation Study on Similarity Measurement. To comprehensively validate the effectiveness of our proposed PCDC-Block, we compare three different similarity measurements $\text{sim}(\cdot,\cdot)$, including (a) traditional inner product in Eq. (3); (b) replacing the PCDC layer in Fig. 5 with a vanilla convolution by inputting the concatenation of $\bar{\bm{q}}$ and $\bar{\bm{k}}$; (c) PCDC-Block in Fig. 5. From the results in Table VI, it is observed that the proposed PCDC-Block has the better capability to capture the relations between query-key pairs for precise similarity calculation and thus promotes the segmentation accuracy. Considering the evident gains over (a), the slight computational cost brought by our method is acceptable. Please refer to SM for visual results.
@@
Ablation Study on Neighbor Selection. Table VII reports the effect of FNS on the segmentation performance achieved by our ReSFU. As seen, without adopting FNS, the model would suffer from a large performance drop, especially for SegFormer with a higher-ratio ×$\times 8$ upsampling process as presented in Fig. 8 (c) (lower). For better clarity, we provide visual comparisons of upsampled features extracted by these two different network structures. From Fig. 14, we can clearly find that for different levels of features, i.e., $\bm{x}$ in Segmenter, $C2$, $C3$, and $C4$ in SegFormer, their upsampled feature maps without FNS always suffer from severe mosaic effects, especially for larger upsampling ratios. Fortunately, this unfavorable phenomenon can be effectively eliminated by adopting our proposed FNS strategy. It is worth mentioning that as explained in Sec. 4.3, FNS is directly implemented based on bilinearly-upsampled features, and it introduces no extra parameters and computational costs.
Table IX: Comparisons on the number of parameters, FLOPs, inference time, and training time per iteration, tested on Segmenter-ViT-S with a 512$\times$512 image on a V100 GPU.
Method Type Params FLOPs Inference Training
Bilinear w/o guidance 22.04M 32.77G 0.012s 0.060s
Deconv 24.40M 35.16G 0.014s 0.062s
PixelShuffle [10] 24.41M 35.16G 0.013s 0.061s
CARAFE [9] 22.21M 32.96G 0.013s 0.061s
DySample [16] 22.09M 32.90G 0.013s 0.061s
Stack-JBU [24] w/ guidance 22.20M 35.32G 0.019s 0.080s
FADE [15] 22.14M 33.92G 0.034s 0.133s
SAPA [2] 22.09M 33.28G 0.013s 0.062s
ReSFU 22.11M 33.88G 0.015s 0.070s
@@
Analysis about Model Complexity. Table IX reports the model complexities of different comparison methods, including the number of parameters, FLOPs, average inference time and training time per iteration based on Segmenter-ViT-S. We can find that 1) Compared to the guidance-free methods, due to the introduction of the HR feature for extra computation, the guidance-based upsampling methods generally cause a certain increase in model complexity and inference/training time; 2) Among the guidance-based feature upsampling methods, our ReSFU still demonstrates stronger competitiveness, which is comparable to SAPA and obviously outperforms other methods, including Stack-JBU and FADE. As seen, our ReSFU performs very competitively in terms of both effectiveness and efficiency. More results are provided in SM.
@@ 5.4.1 Medical Image Segmentation
Performance Comparison. The experiments are executed based on two widely-adopted datasets, i.e., Synapse for multi-organ segmentation, [^3] https://www.synapse.org/#!Synapse:syn3193805/wiki/217789 and the automated cardiac diagnosis challenge (ACDC). [^4] https://www.creatis.insa-lyon.fr/Challenge/acdc/ Specifically, Synapse consists of computed tomography (CT) scans annotated with eight abdominal organs, and ACDC includes magnetic resonance imaging (MRI) scans with three types of annotations. We follow [44] for constructing the training and evaluation datasets. Quantitative evaluation is conducted based on Dice (the higher the better) and the 95-th percentile of the Hausdorff Distance (HD95, the lower the better).
@@ 5.4.2 Instance Segmentation and Panoptic Segmentation
Specifically, for instance segmentation, we employ the classic Mask R-CNN [45] with ResNet50 as the backbone and replace the upsampling stages in FPN with ReSFU, as shown in Fig. 8 (d) (lower). All the results are quantitatively evaluated based on metrics of the average precision (AP) series, including mask AP, AP50, AP75, APS, APM, and APL. For panoptic segmentation, following [16], we choose Panoptic FPN [6] with ResNet50 as the backbone and adjust the three upsamplers in FPN, as presented in Fig. 8 (d) (lower). We report panoptic quality (PQ), PQ on things (PQth), PQ on stuff (PQst), segmentation quality (SQ), and recognition quality (RQ) [6]. The MS COCO dataset [46] is used for training and evaluating the two tasks.
@@
Table XIII: Object detection results under Faster R-CNN-ResNet50 with the default nearest upsampler on MS COCO dataset, and monocular depth estimation results under DepthFormer-Swin-T with the default bilinear upsampler on NYU Depth V2. The best and second-best results are highlighted in bold and underlined, respectively.
Method Object Detection Monocular Depth Estimation
$AP$ $AP_{50}$ $AP_{75}$ $AP_{S}$ $AP_{M}$ $AP_{L}$ $\delta<1.25$()($\uparrow$)( ↑ ) $\delta<1.25^{2}$()($\uparrow$)( ↑ ) $\delta<1.25^{3}$()($\uparrow$)( ↑ ) Abs Rel()($\downarrow$)( ↓ ) RMS()($\downarrow$)( ↓ ) log10()($\downarrow$)( ↓ ) RMS(log)()($\downarrow$)( ↓ ) Sq Rel()($\downarrow$)( ↓ )
Default 37.5 58.2 40.8 21.3 41.1 48.9 0.873 0.978 0.994 0.120 0.402 0.050 0.148 0.071
Stack-JBU [24] 37.5 58.5 40.5 22.1 41.5 47.8 0.874 0.979 0.995 0.118 0.401 0.050 0.148 0.071
CARAFE [9] 38.6 59.9 42.2 23.3 42.2 49.7 0.877 0.978 0.995 0.116 0.397 0.049 0.146 0.069
FADE [15] 38.5 59.6 41.8 23.1 42.2 49.3 0.874 0.978 0.994 0.118 0.399 0.049 0.147 0.071
SAPA [2] 37.8 59.2 40.6 22.4 41.4 49.1 0.870 0.978 0.995 0.117 0.406 0.050 0.149 0.069
DySample [16] 38.6 59.8 42.1 22.5 42.1 50.0 0.872 0.978 0.994 0.119 0.398 0.050 0.148 0.070
ReSFU 38.7 59.9 42.1 23.1 42.3 50.2 0.878 0.979 0.995 0.117 0.395 0.049 0.146 0.070
5.4.3 Object Detection and Monocular Depth Estimation
In addition to segmentation tasks, we further evaluate ReSFU’s potential based on two additional application areas, including object detection and monocular depth estimation. Following the experimental settings adopted in DySample, we conduct experiments of object detection with Faster R-CNN [7] on the MS COCO dataset [46], and conduct experiments of monocular depth estimation with DepthFormer [8] on the NYU Depth V2 dataset [47]. Both Faster R-CNN and DepthFormer contain an FPN neck-based upsampling module as shown in Fig. 8 (d). The corresponding quantitative results are reported in Table XIII. As seen, for these two challenging tasks, almost all methods have very limited room for improvement. However, the overall performance of our ReSFU still surpasses almost all comparison methods in these two tasks, which highlights the favorable universality of our methodological designs.
Table XIV: Semantic segmentation results on ADE20K based on Segmenter and SegFormer with adopting different upsampling manners. The “iterative” upsampling manners for Segmenter and SegFormer are illustrated in Fig. 8 (b)(upper) and Fig. 8 (c)(upper), respectively, while the “direct” upsampler manner is illustrated in the lower row of Fig. 8 (b) and Fig. 8 (c), respectively. The best and second-best results are highlighted in bold and underlined, respectively.
Manner Method Segmenter-ViT-S Segmenter-ViT-L SegFormer-B1 SegFormer-B5
mIoU mAcc bIoU mIoU mAcc bIoU mIoU mAcc bIoU mIoU mAcc bIoU
Iterative CARAFE [9] 46.25 57.63 28.99 51.85 62.96 34.49 42.75 53.96 27.09 50.45 62.29 33.78
FADE [15] 45.71 56.81 28.72 50.07 61.79 33.32 43.09 53.89 28.33 50.43 61.99 34.51
SAPA [2] 45.79 57.36 28.97 51.20 62.46 33.31 42.74 53.39 27.61 49.61 61.11 34.02
DySample [16] 45.78 56.88 28.37 52.04 63.27 34.17 43.48 53.82 27.38 50.14 61.51 33.93
ReSFU 47.02 57.99 30.04 52.53 63.58 35.19 43.84 55.08 28.38 50.55 62.00 34.82
Direct Bilinear 45.75 56.88 27.82 50.96 62.05 33.07 40.97 51.73 24.89 49.13 60.69 32.32
CARAFE [9] 46.10 57.39 28.78 51.73 62.80 34.25 42.65 53.67 27.04 50.22 61.98 33.62
FADE [15] 45.52 56.81 28.13 50.93 62.13 32.60 42.54 53.38 28.11 50.12 61.75 34.10
SAPA [2] 45.59 57.01 28.70 51.04 62.36 33.18 42.41 53.35 27.42 49.55 61.06 33.88
DySample [16] 45.09 55.98 27.73 51.55 63.35 33.66 43.11 53.88 26.53 49.94 61.32 33.62
ReSFU 46.97 57.91 29.96 52.56 63.63 35.27 43.84 54.75 28.06 50.89 62.00 34.40
Figure 19: Visual comparison of upsampled features $\bm{\tilde{x}}$ obtained in an iterative upsampling (Fig. 8 (a) (upper)) and a direct ×$\times 4$ upsampling manners (Fig. 8 (a) (lower)) based on PSPNet-ResNet101.
@@ 5.5 More Discussions on Direct/Iterative Upsampling
Based on the non-hierarchical Segmenter (Fig. 8 (b)) and the hierarchical SegFormer (Fig. 8 (c)), we provide quantitative segmentation results of different comparison methods with adopting different upsampling manners. As reported in Table XIV, for other upsampling methods, they all achieve inferior performance in the direct upsampling manner. In contrast, ReSFU is capable of handling both of these upsampling manners with slight performance fluctuations. Besides, we use PSPNet (Fig. 8 (a)) as an exemplar architecture for a visual demonstration. For the comparing methods, e.g., CARAFE, FADE, SAPA, DySample, and ReSFU, we visually compare their upsampled features $\bm{\tilde{x}}$ obtained in the iterative ×$\times 2$ upsampling manner (Fig. 8 (a) (upper)) and the direct ×$\times 4$ upsampling manner (Fig. 8 (a) (lower)). As displayed in Fig. 19, with the high-ratio direct upsampling, CARAFE, FADE, SAPA, and DySample generate varying degrees of artifacts with more mosaics or blurrier boundaries compared with the iterative upsampling manner. However, our proposed ReSFU achieves better visual effects with richer semantics and clearer boundaries under both upsampling manners. These advantages are essentially attributed to our methodological designs.
As validated, ReSFU achieves similar performance in the direct and iterative upsampling manners, while other comparison methods suffer from performance degradation in the direct manner both numerically and visually. Therefore, in all the experimental comparisons, for previous upsampling methods, we follow their default configuration and adopt them in an iterative manner for fair comparison. Nevertheless, for our ReSFU, we select the direct high-ratio upsampling manner as the default setting in experiments, which can eliminate the requirement for iterative guidance features and facilitate the deployment.
@@ 6 Conclusion and Future Work
In this paper, for the fundamental upsampling design included in almost all current network architectures, we carefully reformulated the current similarity-based feature upsampling framework and thoroughly analyzed the limitations in methodological designs together with experimental visualizations. For each component involved in this pipeline, we meticulously proposed specific optimization designs and correspondingly constructed a refreshed similarity-based feature upsampling framework, called ReSFU. Through comprehensive model verification and ablation studies, we fully validated the role of each component and clearly shown the working mechanism underlying our proposed ReSFU. Based on different types of network architectures, extensive experiments substantiated the superiority of our proposed ReSFU beyond the existing baselines and demonstrated that our ReSFU can be adapted not only to different upsampling structures with flexible guidance manners but also to different applications, including medical image segmentation, instance segmentation, panoptic segmentation, objection detection, and monocular depth estimation. However, for the applications without HR reference images, e.g., image super-resoultion [48] and image generation [49], these guidance-based feature upsampling methods, e.g., FADE, SAPA, and our ReSFU, are difficult to apply. This is beyond our focus in this paper and deserves further investigation in our future work.